OpenStudy (anonymous):

Samantha and Georgia work in different positions at a marketing firm. Samantha's salary is s dollars a month. If Georgia's salary is g dollars a month, what is the best interpretation of the following equation in everyday language? g = s - 0.25s

OpenStudy (texaschic101):

lets just say Samantha's salary is 1000 per month. That means Georgia's salary is : g = 1000 - 0.25(1000) = 1000 - 250 = 750 as you can see, Georgia makes 25% less then Samantha